# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 8.2 Subdirect irredundant subpowers
Proof.
Let $\mathbf{A}$ be a minimal counterexample to the theorem. By Theorem 3.2, $A$ is connected by edges and it suffices to show that each pair of elements connected by an edge is connected by minimal edges. Let $(a,b)$ be an edge in $\mathbf{A}$ and let $\theta$ be a maximal congruence on $\operatorname{Sg}_{\mathbf{A}}(a,b)$ witnessing the edge. The equivalence classes of $\theta$ are connected by minimal edges by the minimality of $\mathbf{A}$. Choose $a^{\prime},b^{\prime}$ such that $(a,a^{\prime}),(b,b^{\prime})\in\theta$ and $\mathbf{B}=\operatorname{Sg}_{\mathbf{A}}(a^{\prime},b^{\prime})$ is inclusion minimal. If $B=A$, then $(a,b)$ is a minimal edge and we are done. Otherwise $a^{\prime}$ and $b^{\prime}$ are connected by minimal edges by the minimality of $\mathbf{A}$, and then so are $a$ and $b$. ∎
@@ 9.1 General facts
Proof.
Let $\mathbf{A}$ be a Taylor algebra and $p>|A|$ be a prime number. By Theorem 3.5 the algebra $\mathbf{A}$ has a cyclic term operation of a prime arity $p$. Consider a family of clones generated by such cyclic operations (the family is finite), and choose a term $t$ which defines a minimal (under inclusion) clone in this family. The algebra $(A;t)$ is clearly a Taylor reduct of $\mathbf{A}$. If $(A;t)$ had a proper Taylor reduct then, by Theorem 3.5, it would have a $p$-ary cyclic term operation. This contradicts the choice of $t$ and shows that $(A;t)$ is a minimal Taylor algebra. ∎
@@
Proof.
Choose a prime number $p>|A|$. The set $B$ together with the restriction of $f$ to $B$, call it $f^{\prime}$, forms a Taylor algebra. By Theorem 3.5 $(B;f^{\prime})$ has a cyclic operation of arity $p$. This operation is defined by a term in $f^{\prime}$, and we let $h$ to be the operation of $\mathbf{A}$ defined by the same term after replacing $f^{\prime}$ by $f$. The set $B$ is closed under $h$ and $h$ is a cyclic operation on $B$. Let $s$ be a cyclic composition of $h$ and a cyclic operation of $\mathbf{A}$ of arity $p$ denoted by $t$ ($t$ exists by Theorem 3.5). The operation $s$ is cyclic and, more importantly, preserves $B$ as $t$ is idempotent and $s$ is cyclic on $B$. Since $\mathbf{A}$ is minimal Taylor, $\operatorname{Clo}(\mathbf{A})=\operatorname{Clo}(A;s)$ and $B$ is a subuniverse of $\mathbf{A}$. ∎
@@
Proof.
For finite powers the claim follows from the definition of a power. Let $\mathbf{A}$ be a minimal Taylor algebra and let $\mathbf{B}$ be a subalgebra or quotient of $\mathbf{A}$. We choose a prime number $p>|A|$ and a $p$-ary cyclic term operation $t$ of $\mathbf{A}$. Using Theorem 3.5 and Proposition 5.2 we find $s\in\operatorname{Clo}(\mathbf{A})$ such that $B$ together with the corresponding term operation $s^{\mathbf{B}}$ of $\mathbf{B}$ is minimal Taylor. Then the cyclic composition $h$ of $t$ and $s$ is a cyclic operation on $\mathbf{A}$ and the corresponding $h^{\mathbf{B}}$ coincides with $s^{\mathbf{B}}$. Since $\mathbf{A}$ is minimal Taylor, we have $\operatorname{Clo}(\mathbf{A})=\operatorname{Clo}(A;h)$ and therefore $\operatorname{Clo}(\mathbf{B})=\operatorname{Clo}(B;h^{\mathbf{B}})=\operatorname{Clo}(B;s^{\mathbf{B}})$, which completes the proof. ∎
@@
Any term operation of a minimal Taylor algebra $\mathbf{A}$ can be obtained by identifying and permuting coordinates (and adding dummy coordinates) of a cyclic term operation of $\mathbf{A}$.
Proof.
Since $\mathbf{A}$ is minimal Taylor, $\operatorname{Clo}(\mathbf{A})=\operatorname{Clo}(A;t)$ for any cyclic operation $t\in\operatorname{Clo}(\mathbf{A})$ (which exists by Theorem 3.5). The claim now follows by noting that the star composition of cyclic operations is a cyclic operation and that, since $t$ is idempotent, any term operation defined by a term in the symbol $t$ can be defined by star composing $t$ multiple times and then permuting and identifying coordinates. ∎
@@ 9.2 Absorption
Suppose $\mathbf{A}$ is a minimal Taylor algebra, $\emptyset\neq B\subsetneq C\subseteq A$, and $\operatorname{Sg}_{\mathbf{A}}(C^{n}\setminus B^{n})\cap B^{n}=\varnothing$ for every $n$. Then for every $f\in\operatorname{Clo}_{n}(\mathbf{A})$ and every essential coordinate $i$ of $f$ we have $f({\mathbf{a}})\notin B$ whenever ${\mathbf{a}}\in C^{n}$ is such that $a_{i}\in C\setminus B$.
Proof.
Any cyclic term operation satisfies the required property (by using the compatibility with $\operatorname{Sg}_{\mathbf{A}}(C^{n}\setminus B^{n})$ on cyclic permutations of ${\mathbf{a}}$) and the property is stable under identifying and permuting coordinates (and introducing dummy ones). The claim now follows from Proposition 9.1. ∎
@@
Proof.
Let $f$ be a witness for $B$ absorbing $\mathbf{A}$ and assume, for a contradiction, that $B$ is not a subuniverse.
Let $\mathbf{A}^{\prime}$ be a reduct of $\mathbf{A}$ with all the operations from $\operatorname{Clo}(\mathbf{A})$ that preserve $B$. Since $\mathbf{A}^{\prime}$ is a proper reduct, it is not a Taylor algebra, and so some quotient of a subalgebra of $\mathbf{A}^{\prime}$ is a two-element algebra such that every operation restricts to a projection (see Proposition 2.2). Let $B_{0}$ and $B_{1}$ be the congruence classes in this quotient, clearly every operation $t$ from $\operatorname{Clo}(\mathbf{A}^{\prime})$ acts like a projection on $\{B_{0},B_{1}\}$. Note that $f$ preserves $B$, therefore it has this property and we assume, without loss of generality, that $f$ acts like the first projection on $\{B_{0},B_{1}\}$.
It follows from the previous paragraph that, for every $n$, the relation $S_{n}=(B_{0}\cup B_{1})^{n}\setminus B_{0}^{n}$ (just like any other relation “built” out of the blocks $B_{0}$ and $B_{1}$) is compatible with every operation from $\operatorname{Clo}(\mathbf{A}^{\prime})$ and is thus pp-definable from $\operatorname{Inv}(\mathbf{A})$ and $B$. Let $T_{n}$ be the relation defined by the same pp-definition with each conjunct $B(x)$ replaced by the void $A(x)$. Since $B$ absorbs $\mathbf{A}^{\prime}$ by $f$, we also know that $S_{n}$ absorbs $T_{n}$ by the same $f$ (see Lemma 7.1)
Suppose that $T_{n}\cap B_{0}^{n}\neq\emptyset$ and choose ${\mathbf{a}}\in T_{n}\cap B_{0}^{n}$ and ${\mathbf{b}}\in B_{1}^{n}$. Then $f({\mathbf{a}},{\mathbf{b}},\dots,{\mathbf{b}})$ belongs to $B_{0}^{n}$ because $f$ acts like the first projection on $\{B_{0},B_{1}\}$, and it also belongs to $S_{n}$ because $S_{n}$ absorbs $T_{n}$ by $f$. This contradiction shows that $T_{n}\cap B_{0}^{n}=\emptyset$ for every $n$. Note that $S_{n}\subseteq T_{n}$ and $T_{n}$ is a subpower of $\mathbf{A}$, hence $\operatorname{Sg}_{\mathbf{A}^{n}}((B_{0}\cup B_{1})^{n}\setminus B_{0}^{n})\cap B_{0}^{n}=\emptyset$ for every $n$. We apply Lemma 9.2 and consider the behaviour of $f$: it preserves $B_{0}\cup B_{1}$ and has at least two essential coordinates (since it is a witness for a nontrivial absorption) and therefore cannot act like a projection on $\{B_{0},B_{1}\}$ — this is a contradiction. ∎
@@
Proof.
We begin with the implications that were discussed in the main part of the paper. The implications from (d) to (c) and from (c) to (b) hold for all algebras. The first one is trivial, and the second one follows immediately from the definitions. The implication from (b) to (a) fails in a trivial clone, but holds in Taylor algebras. To see this we assume (b) and let $t$ be a cyclic term operation of $\mathbf{A}$ of arity $12k+$ 1. Define $f(x,y)=t(x,\dotsc,x,y,\dotsc y)$ where $x$ appears exactly $k+1$ times. It is easy to see that $f$ witnesses the 2-absorption of $B$. Indeed, take any $a\in A,b\in B$ and let ${\mathbf{c}}_{1}=(a,\dotsc,a,b,\dotsc,b)$ with exactly $k+1$ $a$’s. Let ${\mathbf{c}}_{2}$ be a cyclic shift of ${\mathbf{c}}_{1}$ by $k$ positions and ${\mathbf{c}}_{3}$ by $k+1$ positions. Clearly $(t({\mathbf{c}}_{1}),t({\mathbf{c}}_{2}),t({\mathbf{c}}_{3}))\in R$ and, as it is a constant tuple, we conclude that $t({\mathbf{c}}_{1})=f(a,b)\in B$. The case of $f(b,a)$ is similar, so we are done showing that (b) implies (a).
For the implication from (a) to (d) let $g$ be a binary operation witnessing the $2$-absorption of $B$ into $A$ and let $t\in\operatorname{Clo}(\mathbf{A})$ be a cyclic operation of arity $p$. Define $h(x_{1},\dotsc,x_{p})$ as
@@
and note that $h(a_{1},\dotsc,a_{p})\in B$ whenever at least one of the $a_{i}$ is in $B$. The same property holds for the cyclic composition, call it $s$, of $t$ and $h$. To see this, take a $p$-tuple ${\mathbf{a}}$ which, in some position, has an element from $B$. In the subterms of $s$ the operation $h$ is applied to cyclic shifts of ${\mathbf{a}}$, and every time the result is in $B$. Then $t$ is applied to elements from $B$ and, by Theorem 5.5, the result is in $B$ as well. The operation $s$ is cyclic and therefore generates the whole clone of $\mathbf{A}$. Thus if any variable appearing in a term $f$ build from $s$ (in particular, a variable corresponding to an essential coordinate of $f^{\mathbf{A}}$) is an element of $B$, then the whole term evaluates to an element of $B$. ∎
@@
Proof.
For (1) consider the result of applying an operation $f\in\operatorname{Clo}(\mathbf{A})$ to a tuple ${\mathbf{a}}$. If $a_{i}\in C$ for all of the essential coordinates $i$ of $f$, then the result is in $C$ (as $C\leq\mathbf{A}$), and if $a_{i}\in B$ for an essential $i$, then the result is in $B$ by item (d) in Theorem 5.7.
The argument for (2.a) is similar. From (1) we know that $B\cup C$ is a subuniverse. For the absorption note that if a tuple ${\mathbf{a}}$ has all but one entry in $B\cup C$, then $f({\mathbf{a}})\in B$ in the case where $a_{i}\in B$ for some essential $i$, or $f({\mathbf{a}})\in C$ in the other case (as $C\trianglelefteq\mathbf{A}$ by $f$ and the inessential coordinates can be substituted with elements of $C$). For (2.b) first note that $f$ has at least two essential coordinates. Then $B\cap C\trianglelefteq\mathbf{A}$ follows again from item (d) in Theorem 5.7, and $B\cap C$ is nonempty since it contains $f(c,\dotsc,c,b,c,\dotsc,c)$ for any $b\in B$ and $c\in C$ (where $b$ is at an essential coordinate).
To prove (3) note that, by strong projectivity, both absorptions $\mathbf{C}\trianglelefteq_{2}\mathbf{B}\trianglelefteq_{2}\mathbf{A}$ can be witnessed by the same operation $f$ (and, in fact, any binary term operation of $\mathbf{A}$ such that both coordinates are essential can be taken as a witness) and then $f(f(f(x,y),x),f(f(y,x),y))$ witnesses $\mathbf{C}\trianglelefteq_{2}\mathbf{A}$
The first part of item (4) follows from (2.b) as, again, all 2-absorptions have a common witness. The second part follows from (3). ∎
@@
Proof.
Let $f,f_{1},\dotsc,f_{n}$ and $g,g_{1},\dotsc,g_{n}$ be as in the statement of the proposition. Take any ${\mathbf{a}}\in A^{k}$ and let ${\mathbf{x}}$ be its characteristic $k$-tuple (with respect to $B$). Assume that $g(g_{1},\dotsc,g_{n})({\mathbf{x}})=1$ and note that if $g_{i}({\mathbf{x}})=1$ then $f_{i}({\mathbf{a}})\in B$. Therefore the characteristic $n$-tuple of $(f_{1}({\mathbf{a}}),\dotsc,f_{n}({\mathbf{a}}))$ is above the tuple $(g_{1}({\mathbf{x}}),\dots,g_{n}({\mathbf{x}}))$ and since $g$ applied to the latter tuple outputs $1$, it outputs $1$ on the former as well. By the definition of $\rightsquigarrow_{B}$, this implies that $f(f_{1}({\mathbf{a}}),\dotsc,f_{n}({\mathbf{a}}))\in B$, and the proposition is proved. ∎
Note that the previous proposition is especially useful when the algebra $\mathbf{A}$ is generated by a single operation. More formally if $B\subseteq A$ and $f\rightsquigarrow_{B}g$, then any $f^{\prime}\in\operatorname{Clo}(A;f)$ is $\rightsquigarrow_{B}$ related to some $g^{\prime}\in\operatorname{Clo}(\{0,1\};g)$. Indeed, for any fixed $k$, it suffices to start with $\operatorname{proj}^{k}_{i}\mbox{(on $A$)}\rightsquigarrow_{B}\operatorname{proj}^{k}_{i}$ (on $\{0,1\}$), for all $i\leq k$, and apply the previous proposition as many times as needed. This is in fact exactly how we prove “(a) implies (d)” in Theorem 5.10 – the proof of this implication is extracted to the following theorem.
@@
Let $\mathbf{A}$ be a minimal Taylor algebra. There exists an arity preserving map $f\mapsto f^{\prime}$ from $\operatorname{Clo}(\mathbf{A})$ to $\operatorname{Clo}(\{0,1\};\operatorname{maj})$ such that:
  • if a set $B$ $3$-absorbs $A$ then $f\rightsquigarrow_{B}f^{\prime}$ for every $f$ in $\operatorname{Clo}(\mathbf{A})$, and
@@
  • for every prime $p>|A|$ there is a cyclic term of $\mathbf{A}$ of arity $p$ which is mapped to $\operatorname{maj}_{p}$.
Proof.
@@
For any prime $p>|A|$, we have a $p$-ary cyclic operation, say $t$, in $\mathbf{A}$. If $t$ satisfies $t\rightsquigarrow_{B}\operatorname{maj}_{p}$ then one can generate, as in discussion after Proposition 9.4, a subdirect relation $\rightsquigarrow_{B}$ between $\operatorname{Clo}(\mathbf{A})$ and $\operatorname{Clo}(\{0,1\};\operatorname{maj})$. Once this is done, we see that for any $f\in\operatorname{Clo}(\mathbf{A})$ we can choose any operation $\rightsquigarrow_{B}$-related to $f$ and fix it to be $f^{\prime}$. Such a map satisfies the conclusion of the theorem, for the set $B$ in question. If the initial condition, i.e. $t\rightsquigarrow_{B}\operatorname{maj}_{p}$, held for a number of $B$’s then the conclusion will hold for all of them. We will prove the theorem by induction on the size of the set of $B$’s.
@@
Given a set $\mathcal{B}$ of 3-absorbing subsets of $A$ and a cyclic operation $t\in\operatorname{Clo}_{p}(\mathbf{A})$, such that $t\rightsquigarrow_{B}\operatorname{maj}_{p}$ for every $B\in\mathcal{B}$, we will find another cyclic term operation $s$ which will still work for any $B\in\mathcal{B}$ and, additionally, for a new 3-absorbing subset $C$ in $\mathbf{A}$. The claim will then follow by induction since we can start with the empty $\mathcal{B}$ and any $p$-ary cyclic term operation of $\mathbf{A}$ (in such a case we can choose an arbitrary map $f\mapsto f^{\prime}$ sending $t$ to $\operatorname{maj}_{p}$).
@@
Let $f\mapsto f^{\prime}$ be a map compatible with $\rightsquigarrow_{B}$ for all $B\in\mathcal{B}$ and such that $t\mapsto\operatorname{maj}_{p}$. Let $f$ be a witness for $C$ 3-absorbing $\mathbf{A}$, and let $f^{\prime}$ be the image of $f$. Finally, let $h$ be the $p$-ary term operation of $\mathbf{A}$ defined from $f$ by the same term as a term defining $\operatorname{maj}_{p}$ from $\operatorname{maj}$ (the latter term exists since $\operatorname{Clo}(\{0,1\},\operatorname{maj})$ is the clone of monotone selfdual operations – see the discussion in Section 5) and let $s$ be the cyclic composition of $t$ and $h$. Our aim is to verify that $s\rightsquigarrow_{B}\operatorname{maj}_{p}$ for every $B\in\mathcal{B}\cup\{C\}$.
@@
Observe that $f\rightsquigarrow_{C}\operatorname{maj}$ by the definition of absorption, therefore $h\rightsquigarrow_{C}\operatorname{maj}_{p}$ by the definition of $h$ and Proposition 9.4. Further $s\rightsquigarrow_{C}\operatorname{maj}_{p}$ by the definition of $s$ and the fact that $C\leq\mathbf{A}$ (which follows from Theorem 5.5). We thus further concentrate on the case $B\in\mathcal{B}$. If $f^{\prime}$ (the image of $f$ from the previous paragraph) is $\operatorname{maj}$, then $h\rightsquigarrow_{B}\operatorname{maj}_{p}$ and further $s\rightsquigarrow_{B}\operatorname{maj}_{p}$ for all $B\in\mathcal{B}$ by the same reasoning, so this case is done.

### v6

@@ 8.2 Subdirect irredundant subpowers
Proof 8.33.
Let be a minimal counterexample to the theorem. By Theorem 3.2, $A$ is connected by edges and it suffices to show that each pair of elements connected by an edge is connected by minimal edges. Let $(a,b)$ be an edge in and let $\theta$ be a maximal congruence on $\operatorname{Sg}(a,b)$ witnessing the edge. The equivalence classes of $\theta$ are connected by minimal edges by the minimality of . Choose $a^{\prime},b^{\prime}$ such that $(a,a^{\prime}),(b,b^{\prime})\in\theta$ and is inclusion minimal. If $B=A$, then $(a,b)$ is a minimal edge and we are done. Otherwise $a^{\prime}$ and $b^{\prime}$ are connected by minimal edges by the minimality of , and then so are $a$ and $b$.
@@ 9.1 General facts
Proof 9.1.
Let be a Taylor algebra and $p>|A|$ be a prime number. By Theorem 3.5 the algebra  has a cyclic term operation of a prime arity $p$. Consider a family of clones generated by such cyclic operations (the family is finite), and choose a term $t$ which defines a minimal (under inclusion) clone in this family. The algebra $(A;t)$ is clearly a Taylor reduct of . If $(A;t)$ had a proper Taylor reduct then, by Theorem 3.5, it would have a $p$-ary cyclic term operation. This contradicts the choice of $t$ and shows that $(A;t)$ is a minimal Taylor algebra.
@@
Proof 9.2.
Choose a prime number $p>|A|$. The set $B$ together with the restriction of $f$ to $B$, call it $f^{\prime}$, forms a Taylor algebra. By Theorem 3.5 $(B;f^{\prime})$ has a cyclic operation of arity $p$. This operation is defined by a term in $f^{\prime}$, and we let $h$ to be the operation of defined by the same term after replacing $f^{\prime}$ by $f$. The set $B$ is closed under $h$ and $h$ is a cyclic operation on $B$. Let $s$ be a cyclic composition of $h$ and a cyclic operation of of arity $p$ denoted by $t$ ($t$ exists by Theorem 3.5). The operation $s$ is cyclic and, more importantly, preserves $B$ as $t$ is idempotent and $s$ is cyclic on $B$. Since is minimal Taylor, $\operatorname{Clo}($ and $B$ is a subuniverse of .
@@
Proof 9.3.
For finite powers the claim follows from the definition of a power. Let be a minimal Taylor algebra and let be a subalgebra or quotient of . We choose a prime number $p>|A|$ and a $p$-ary cyclic term operation $t$ of . Using Theorem 3.5 and Proposition 5.2 we find $s\in\operatorname{Clo}($ such that $B$ together with the corresponding term operation $s$ of is minimal Taylor. Then the cyclic composition $h$ of $t$ and $s$ is a cyclic operation on and the corresponding $h$ coincides with $s$. Since is minimal Taylor, we have $\operatorname{Clo}($ and therefore $\operatorname{Clo}($, which completes the proof.
@@
Any term operation of a minimal Taylor algebra can be obtained by identifying and permuting coordinates (and adding dummy coordinates) of a cyclic term operation of .
Proof 9.5.
Since is minimal Taylor, $\operatorname{Clo}($ for any cyclic operation $(t\in\operatorname{Clo}($which exists by Theorem 3.5). The claim now follows by noting that the star composition of cyclic operations is a cyclic operation and that, since $t$ is idempotent, any term operation defined by a term in the symbol $t$ can be defined by star composing $t$ multiple times and then permuting and identifying coordinates.
@@ 9.2 Absorption
Suppose is a minimal Taylor algebra, $\emptyset\neq B\subsetneq C\subseteq A$, and $\operatorname{Sg}_{\mathbf{A}}(C^{n}\setminus B^{n})\cap B^{n}=\varnothing$ for every $n$. Then for every $f\in\operatorname{Clo}_{n}($ and every essential coordinate $i$ of $f$ we have $f({\mathbf{a}})\notin B$ whenever ${\mathbf{a}}\in C^{n}$ is such that $a_{i}\in C\setminus B$.
Proof 9.7.
Any cyclic term operation satisfies the required property (by using the compatibility with $\operatorname{Sg}_{\mathbf{A}}(C^{n}\setminus B^{n})$ on cyclic permutations of ${\mathbf{a}}$) and the property is stable under identifying and permuting coordinates (and introducing dummy ones). The claim now follows from Proposition 9.4.
@@
Proof 9.8.
Let $f$ be a witness for $B$ absorbing and assume, for a contradiction, that $B$ is not a subuniverse.
Let be a reduct of with all the operations from $\operatorname{Clo}($ that preserve $B$. Since is a proper reduct, it is not a Taylor algebra, and so some quotient of a subalgebra of is a two-element algebra such that every operation restricts to a projection (see Proposition 2.2). Let $B_{0}$ and $B_{1}$ be the congruence classes in this quotient, clearly every operation $t$ from $\operatorname{Clo}($ acts like a projection on $\{B_{0},B_{1}\}$. Note that $f$ preserves $B$, therefore it has this property and we assume, without loss of generality, that $f$ acts like the first projection on $\{B_{0},B_{1}\}$.
It follows from the previous paragraph that, for every $n$, the relation $S_{n}=(B_{0}\cup B_{1})^{n}\setminus B_{0}^{n}$ (just like any other relation “built” out of the blocks $B_{0}$ and $B_{1}$) is compatible with every operation from $\operatorname{Clo}($ and is thus pp-definable from $\operatorname{Inv}($ and $B$. Let $T_{n}$ be the relation defined by the same pp-definition with each conjunct $B(x)$ replaced by the void $A(x)$. Since $B$ absorbs by $f$, we also know that $S_{n}$ absorbs $T_{n}$ by the same $f$ (see Lemma 7.1)
Suppose that $T_{n}\cap B_{0}^{n}\neq\emptyset$ and choose ${\mathbf{a}}\in T_{n}\cap B_{0}^{n}$ and ${\mathbf{b}}\in B_{1}^{n}$. Then $f({\mathbf{a}},{\mathbf{b}},\dots,{\mathbf{b}})$ belongs to $B_{0}^{n}$ because $f$ acts like the first projection on $\{B_{0},B_{1}\}$, and it also belongs to $S_{n}$ because $S_{n}$ absorbs $T_{n}$ by $f$. This contradiction shows that $T_{n}\cap B_{0}^{n}=\emptyset$ for every $n$. Note that $S_{n}\subseteq T_{n}$ and $T_{n}$ is a subpower of $\mathbf{A}$, hence $\operatorname{Sg}_{\mathbf{A}^{n}}((B_{0}\cup B_{1})^{n}\setminus B_{0}^{n})\cap B_{0}^{n}=\emptyset$ for every $n$. We apply Lemma 9.6 and consider the behaviour of $f$: it preserves $B_{0}\cup B_{1}$ and has at least two essential coordinates (since it is a witness for a nontrivial absorption) and therefore cannot act like a projection on $\{B_{0},B_{1}\}$ — this is a contradiction.
@@
Proof 9.9.
We begin with the implications that were discussed in the main part of the paper. The implications from (d) to (c) and from (c) to (b) hold for all algebras. The first one is trivial, and the second one follows immediately from the definitions. The implication from (b) to (a) fails in a trivial clone, but holds in Taylor algebras. To see this we assume (b) and let $t$ be a cyclic term operation of of arity $12k+$ 1. Define $f(x,y)=t(x,\dotsc,x,y,\dotsc y)$ where $x$ appears exactly $k+1$ times. It is easy to see that $f$ witnesses the 2-absorption of $B$. Indeed, take any $a\in A,b\in B$ and let ${\mathbf{c}}_{1}=(a,\dotsc,a,b,\dotsc,b)$ with exactly $k+1$ $a$’s. Let ${\mathbf{c}}_{2}$ be a cyclic shift of ${\mathbf{c}}_{1}$ by $k$ positions and ${\mathbf{c}}_{3}$ by $k+1$ positions. Clearly $(t({\mathbf{c}}_{1}),t({\mathbf{c}}_{2}),t({\mathbf{c}}_{3}))\in R$ and, as it is a constant tuple, we conclude that $t({\mathbf{c}}_{1})=f(a,b)\in B$. The case of $f(b,a)$ is similar, so we are done showing that (b) implies (a).
For the implication from (a) to (d) let $g$ be a binary operation witnessing the $2$-absorption of $B$ into $A$ and let $t\in\operatorname{Clo}($ be a cyclic operation of arity $p$. Define $h(x_{1},\dotsc,x_{p})$ as
@@
and note that $h(a_{1},\dotsc,a_{p})\in B$ whenever at least one of the $a_{i}$ is in $B$. The same property holds for the cyclic composition, call it $s$, of $t$ and $h$. To see this, take a $p$-tuple ${\mathbf{a}}$ which, in some position, has an element from $B$. In the subterms of $s$ the operation $h$ is applied to cyclic shifts of ${\mathbf{a}}$, and every time the result is in $B$. Then $t$ is applied to elements from $B$ and, by Theorem 5.5, the result is in $B$ as well. The operation $s$ is cyclic and therefore generates the whole clone of . Thus if any variable appearing in a term $f$ build from $s$ (in particular, a variable corresponding to an essential coordinate of $f$) is an element of $B$, then the whole term evaluates to an element of $B$.
@@
Proof 9.10.
For (1) consider the result of applying an operation $f\in\operatorname{Clo}($ to a tuple ${\mathbf{a}}$. If $a_{i}\in C$ for all of the essential coordinates $i$ of $f$, then the result is in $C$ (as $C\leq$), and if $a_{i}\in B$ for an essential $i$, then the result is in $B$ by item (d) in Theorem 5.7.
The argument for (2.a) is similar. From (1) we know that $B\cup C$ is a subuniverse. For the absorption note that if a tuple ${\mathbf{a}}$ has all but one entry in $B\cup C$, then $f({\mathbf{a}})\in B$ in the case where $a_{i}\in B$ for some essential $i$, or $f({\mathbf{a}})\in C$ in the other case (as $C\trianglelefteq$ by $f$ and the inessential coordinates can be substituted with elements of $C$). For (2.b) first note that $f$ has at least two essential coordinates. Then $B\cap C\trianglelefteq$ follows again from item (d) in Theorem 5.7, and $B\cap C$ is nonempty since it contains $f(c,\dotsc,c,b,c,\dotsc,c)$ for any $b\in B$ and $c\in C$ (where $b$ is at an essential coordinate).
To prove (3) note that, by strong projectivity, both absorptions can be witnessed by the same operation $f$ (and, in fact, any binary term operation of such that both coordinates are essential can be taken as a witness) and then $f(f(f(x,y),x),f(f(y,x),y))$ witnesses
The first part of item (4) follows from (2.b) as, again, all 2-absorptions have a common witness. The second part follows from (3).
@@
Proof 9.13.
Let $f,f_{1},\dotsc,f_{n}$ and $g,g_{1},\dotsc,g_{n}$ be as in the statement of the proposition. Take any ${\mathbf{a}}\in A^{k}$ and let ${\mathbf{x}}$ be its characteristic $k$-tuple (with respect to $B$). Assume that $g(g_{1},\dotsc,g_{n})({\mathbf{x}})=1$ and note that if $g_{i}({\mathbf{x}})=1$ then $f_{i}({\mathbf{a}})\in B$. Therefore the characteristic $n$-tuple of $(f_{1}({\mathbf{a}}),\dotsc,f_{n}({\mathbf{a}}))$ is above the tuple $(g_{1}({\mathbf{x}}),\dots,g_{n}({\mathbf{x}}))$ and since $g$ applied to the latter tuple outputs $1$, it outputs $1$ on the former as well. By the definition of $\rightsquigarrow_{B}$, this implies that $f(f_{1}({\mathbf{a}}),\dotsc,f_{n}({\mathbf{a}}))\in B$, and the proposition is proved.
Note that the previous proposition is especially useful when the algebra is generated by a single operation. More formally if $B\subseteq A$ and $f\rightsquigarrow_{B}g$, then any $f^{\prime}\in\operatorname{Clo}(A;f)$ is $\rightsquigarrow_{B}$ related to some $g^{\prime}\in\operatorname{Clo}(\{0,1\};g)$. Indeed, for any fixed $k$, it suffices to start with $\operatorname{proj}^{k}_{i}\mbox{(on $A$)}\rightsquigarrow_{B}\operatorname{proj}^{k}_{i}$ (on $\{0,1\}$), for all $i\leq k$, and apply the previous proposition as many times as needed. This is in fact exactly how we prove “(a) implies (d)” in Theorem 5.10 – the proof of this implication is extracted to the following theorem.
@@
Let be a minimal Taylor algebra. There exists an arity preserving map $f\mapsto f^{\prime}$ from $\operatorname{Clo}($ to $\operatorname{Clo}(\{0,1\};\operatorname{maj})$ such that:
  • if a set $B$ $3$-absorbs $A$ then $f\rightsquigarrow_{B}f^{\prime}$ for every $f$ in $\operatorname{Clo}($, and
@@
  • for every prime $p>|A|$ there is a cyclic term of of arity $p$ which is mapped to $\operatorname{maj}_{p}$.
Proof 9.15.
@@
For any prime $p>|A|$, we have a $p$-ary cyclic operation, say $t$, in . If $t$ satisfies $t\rightsquigarrow_{B}\operatorname{maj}_{p}$ then one can generate, as in discussion after Proposition 9.12, a subdirect relation $\rightsquigarrow_{B}$ between $\operatorname{Clo}($ and $\operatorname{Clo}(\{0,1\};\operatorname{maj})$. Once this is done, we see that for any $f\in\operatorname{Clo}($ we can choose any operation $\rightsquigarrow_{B}$-related to $f$ and fix it to be $f^{\prime}$. Such a map satisfies the conclusion of the theorem, for the set $B$ in question. If the initial condition, i.e. $t\rightsquigarrow_{B}\operatorname{maj}_{p}$, held for a number of $B$’s then the conclusion will hold for all of them. We will prove the theorem by induction on the size of the set of $B$’s.
@@
Given a set $\mathcal{B}$ of 3-absorbing subsets of $A$ and a cyclic operation $t\in\operatorname{Clo}_{p}($, such that $t\rightsquigarrow_{B}\operatorname{maj}_{p}$ for every $B\in\mathcal{B}$, we will find another cyclic term operation $s$ which will still work for any $B\in\mathcal{B}$ and, additionally, for a new 3-absorbing subset $C$ in . The claim will then follow by induction since we can start with the empty $\mathcal{B}$ and any $p$-ary cyclic term operation of  (in such a case we can choose an arbitrary map $f\mapsto f^{\prime}$ sending $t$ to $\operatorname{maj}_{p}$).
@@
Let $f\mapsto f^{\prime}$ be a map compatible with $\rightsquigarrow_{B}$ for all $B\in\mathcal{B}$ and such that $t\mapsto\operatorname{maj}_{p}$. Let $f$ be a witness for $C$ 3-absorbing , and let $f^{\prime}$ be the image of $f$. Finally, let $h$ be the $p$-ary term operation of defined from $f$ by the same term as a term defining $\operatorname{maj}_{p}$ from $\operatorname{maj}$ (the latter term exists since $\operatorname{Clo}(\{0,1\},\operatorname{maj})$ is the clone of monotone selfdual operations – see the discussion in Section 5) and let $s$ be the cyclic composition of $t$ and $h$. Our aim is to verify that $s\rightsquigarrow_{B}\operatorname{maj}_{p}$ for every $B\in\mathcal{B}\cup\{C\}$.
@@
Observe that $f\rightsquigarrow_{C}\operatorname{maj}$ by the definition of absorption, therefore $h\rightsquigarrow_{C}\operatorname{maj}_{p}$ by the definition of $h$ and Proposition 9.12. Further $s\rightsquigarrow_{C}\operatorname{maj}_{p}$ by the definition of $s$ and the fact that $C\leq$ (which follows from Theorem 5.5). We thus further concentrate on the case $B\in\mathcal{B}$. If $f^{\prime}$ (the image of $f$ from the previous paragraph) is $\operatorname{maj}$, then $h\rightsquigarrow_{B}\operatorname{maj}_{p}$ and further $s\rightsquigarrow_{B}\operatorname{maj}_{p}$ for all $B\in\mathcal{B}$ by the same reasoning, so this case is done.